# Bayesian Structure Learning in Undirected Gaussian Graphical Models: Literature Review with Empirical Comparison

### v2

@@ 4 Empirical Comparison
The simulation study in this section contributes to the literature in three ways. First, it includes large-scale instances, i.e., instances with a thousand variables. Until now, Bayesian algorithms are only compared on 250 variables or less, an exception being the K-LRD algorithm by Chandra et al. (2022). Second, we report information on how long time it takes algorithms to achieve good results. This information is missing in Bayesian literature. Third, and most importantly, never before has such a variety of Bayesian algorithms been compared in a single simulation study. All our results can be reproduced by the specified scripts on our GitHub page [^1] https://anonymous.4open.science/r/Review-paper-Bayesian-Structure-Learning-in-GGMs-6658. We discuss the simulation setup in Subsection 4.1 and the results in Subsection 4.2.
@@ 4.1 Simulation Setup
Selected algorithms: In Bayesian structure learning literature, new algorithms tend to be compared to frequentist algorithms, or Bayesian algorithms of similar nature, e.g., K-algorithms are compared to K-algorithms, ECM algorithms to ECM algorithms, and so on. Comparisons between dissimilar Bayesian algorithms are often outdated or missing completely. To illustrate this, we group Bayesian methods into seven categories. Table 5 shows that comparisons between different categories are largely missing in the literature.
@@
Frequentist missing missing 2015 2015 missing 2024 2024 2018 2020
@@
In this section, we complete most of this table by comparing seven Bayesian algorithms (RJ - WWA, BD-A, SS, G-MPLBD, K-Horseshoe, ECM-BAGUS, and H-BGGM) and one frequentist algorithm (glasso). For this simulation study, Bayesian algorithms are chosen based on three criteria: i) the algorithm demonstrates outstanding performance in its category, ii) there is publicly accessible working code for the algorithm, and iii) the algorithm is documented in a published article. No regression algorithm is selected because there is no code available. As a prior for the graph we sample the edges from independent and identically distributed Bernoulli distributions. The resulting prior becomes
@@
where $\delta$ is the prior sparsity of the graph. In literature this prior sparsity is often set to either the uninformative value $\delta=0.5$ (Gan et al., 2019) or to a function that decreases with $p$, for example $2/(p-1)$ (Van den Boom et al., 2022a). In this simulation study, we opt for a middle ground and set $\delta=0.2$. Supplementary materials S1 and S2 contain the source and all other settings of the algorithms.
@@
We also measure the computational cost of each algorithm. We first run each Bayesian algorithm until visual inspection shows convergence and denote with $P$ the final edge inclusion matrix. The computational cost $T$ of an algorithm is then defined as the time it takes the algorithm to produce an AUC within $\epsilon=0.01$ of its final AUC. That is,
@@
The value of $\delta$ is chosen so that the resulting edge density in the graphs is $10\%$ for $p=10$, $1\%$ and $10\%$ for $p=100$, and $0.1\%$ for $p=1000$. For every generated graph $G^{*}$, the corresponding precision matrix $\mathbf{K}^{*}$ is then sampled from the $G$-Wishart distribution. That is $\mathbf{K}^{*}\sim W_{G^{*}}(3,\mathbf{I})$. For $p\in\{10,100,1000\}$, each $n$, each graph type, and each density, we repeat the process 16 times which is generally enough to reach low ($<0.025$) standard errors on AUC, $Pr^{+}$ and $Pr^{-}$. For each generated $G^{*}$ and $\mathbf{K}^{*}$, we then sample $n$ observations from the $p-$variate normal distribution with covariance matrix $\mathbf{\Sigma}^{*}=\mathbf{K}^{*{-1}}$ and mean zero.
@@ 4.2 Results
The results on the computational costs $T$, AUC, $Pr^{-}$ and $Pr^{+}$, and are shown, respectively, in Tables 6, 7, 8, and 9. If there is one conclusion to be drawn from these results, then it is this: the idea that Bayesian algorithms lack the computational advantage of their frequentist counterparts is outdated. Whether with 10, 100, or 1000 variables, one or more Bayesian algorithms can estimate the full posterior distribution or any function thereof at the same time that the glasso algorithm can provide a point estimate. Moreover, there is no evidence of the glasso algorithm outperforming Bayesian algorithms on AUC, $Pr^{+}$, or $Pr^{-}$ either. It underlines the staggering progress Bayesian algorithms have made over the last decade. This progress is illustrated too in Figure 1, which shows the convergence of performance metrics for each algorithm on a large scale instance $(p=1000)$: the AUC, $Pr^{+}$ and $Pr^{-}$ values of the G-MPLBD, SS-O, and K-Horseshoe algorithms converge in approximately the same time as the glasso algorithm.
@@
In terms of the AUC, most algorithms perform similarly. Noteworthy is the overall underperformance of the ECM-BAGUS algorithm and the underperformance of the glasso algorithm in medium-sized problems ($p=100$) with a high edge density ($10$%). The $Pr^{-}$ values are comparable across most algorithms too with estimates close to the desired value of zero. Only the ECM-BAGUS and BGGM algorithms perform significantly worse, the latter even assigning in some cases an average edge inclusion probability of $0.88$ to edges that are not in the true graph. Looking at the $Pr^{+}$ values, we see more variability between algorithms. For instances with a high number of observations, the K-Horseshoe algorithm performs best. For large-scale graphs, the K-horseshoe algorithm outperforms the others too. The ECM-BAGUS algorithm performs slightly worse than others in most instances. The $Pr^{+}$ values of the BGGM algorithm seem close to optimal but are only marginally more than their $Pr^{-}$ values.
@@
The results also reveal that it is not so much the structure of the graph (random or cluster) as the density that influences the performance of algorithms. More observations lead to better results. Generally, $n=350\log p$ observations are sufficient to yield good results.
@@
Figure 1: AUC (topleft), $Pr^{+}$ (topright) and $Pr^{-}$ (bottom) over time. The figures show averages over 16 replications on an instance with $p=1000$, $n=1050$ on the random graph with density $0.1\%$.
p graph density n glasso RJ-WWA BD-A SS ECM-BAGUS G-MPLBD K-Horseshoe H-BGGM
10 random 10 % 20 2 (0) 1 (1) 0 (1) 0 (0) 1 (0) 0 (0) 0 (0) 0 (0)
350 2 (0) 0 (1) 0 (0) 0 (0) 1 (0) 0 (0) 0 (1) 0 (0)
cluster 10 % 20 2 (0) 1 (1) 0 (0) 0 (0) 1 (0) 0 (0) 0 (1) 0 (0)
350 2 (0) 0 (0) 0 (0) 0 (0) 1 (0) 0 (0) 0 (0) 0 (0)
100 random 1 % 40 3 (0) 3623 (4183) 5313 (4190) 5 (6) 84 (9) 4 (4) 49 (56) -
700 3 (0) 147 (77) 1835 (2934) 19 (22) 76 (18) 2 (2) 3 (1) 11 (0)
10 % 40 5 (1) 2794 (1513) 9018 (3161) 10 (5) 252 (80) 6 (4) 32 (22) -
700 5 (1) 746 (441) 6093 (2319) 12 (5) 482 (164) 8 (3) 18 (13) 10 (0)
cluster 1 % 40 3 (0) 620 (402) 4496 (3730) 13 (16) 83 (6) 4 (3) 11 (13) -
700 3 (0) 159 (130) 3013 (4124) 19 (17) 70 (12) 1 (2) 5 (3) 11 (0)
10 % 40 4 (0) 3489 (1499) 10368 (2727) 13 (7) 186 (50) 5 (3) 17 (18) -
700 3 (0) 1569 (759) 6465 (3461) 10 (7) 349 (127) 6 (3) 10 (5) 11 (0)
1000 random 0.1 % 60 606 (54) - - 9063 (1538) 15057 (1873) 624 (19) 24064 (417) -
1050 354 (55) - - 4007 (1519) 8907 (1158) 492 (163) 12300 (200) -
cluster 0.1 % 60 672 (149) - - 8510 (1513) 15710 (1214) 623 (35) 7451 (7762) -
1050 366 (84) - - 3823 (1861) 8224 (1539) 481 (198) 12389 (128) -
@@
Table 6: Computational cost $T$ (24) in seconds of the algorithms for different instances. The values are averages and standard deviations (between brackets) over 16 replications. The “-” entry indicates that an algorithm did not produce results on that instance.
p graph density n glasso RJ-WWA BD-A SS ECM-BAGUS G-MPLBD K-Horseshoe H-BGGM
10 random 10 % 20 0.7 (0.24) 0.72 (0.24) 0.73 (0.21) 0.73 (0.22) 0.7 (0.21) 0.72 (0.22) 0.7 (0.23) 0.67 (0.24)
350 0.94 (0.08) 0.95 (0.08) 0.95 (0.08) 0.93 (0.09) 0.91 (0.1) 0.95 (0.07) 0.94 (0.06) 0.93 (0.08)
cluster 10 % 20 0.83 (0.14) 0.82 (0.16) 0.82 (0.16) 0.81 (0.14) 0.79 (0.13) 0.82 (0.13) 0.81 (0.17) 0.75 (0.15)
350 0.93 (0.08) 0.93 (0.1) 0.93 (0.1) 0.91 (0.14) 0.92 (0.1) 0.94 (0.1) 0.92 (0.11) 0.92 (0.12)
100 random 1 % 40 0.83 (0.06) 0.83 (0.07) 0.81 (0.07) 0.84 (0.06) 0.76 (0.06) 0.81 (0.07) 0.84 (0.07) -
700 0.96 (0.03) 0.96 (0.03) 0.96 (0.03) 0.94 (0.02) 0.94 (0.03) 0.95 (0.03) 0.96 (0.02) 0.95 (0.02)
10 % 40 0.67 (0.03) 0.7 (0.01) 0.68 (0.02) 0.71 (0.02) 0.59 (0.02) 0.67 (0.02) 0.71 (0.02) -
700 0.78 (0.06) 0.92 (0.01) 0.91 (0.01) 0.89 (0.01) 0.8 (0.07) 0.9 (0.01) 0.92 (0.01) 0.87 (0.01)
cluster 1 % 40 0.83 (0.04) 0.85 (0.04) 0.83 (0.05) 0.83 (0.04) 0.78 (0.04) 0.81 (0.04) 0.85 (0.05) -
700 0.95 (0.02) 0.96 (0.02) 0.96 (0.02) 0.94 (0.02) 0.94 (0.02) 0.94 (0.02) 0.95 (0.03) 0.94 (0.02)
10 % 40 0.71 (0.02) 0.77 (0.02) 0.72 (0.02) 0.75 (0.02) 0.61 (0.02) 0.7 (0.02) 0.73 (0.01) -
700 0.84 (0.02) 0.94 (0.01) 0.93 (0.01) 0.91 (0.01) 0.84 (0.04) 0.92 (0.01) 0.93 (0.01) 0.87 (0.01)
1000 random 0.1 % 60 0.86 (0.03) - - 0.84 (0.03) 0.84 (0.03) 0.72 (0.04) 0.84 (0.03) -
1050 0.97 (0.01) - - 0.95 (0.01) 0.96 (0.01) 0.95 (0.01) 0.96 (0.01) -
cluster 0.1 % 60 0.84 (0.04) - - 0.82 (0.05) 0.84 (0.03) 0.7 (0.05) 0.84 (0.04) -
1050 0.96 (0.01) - - 0.95 (0.02) 0.95 (0.02) 0.95 (0.01) 0.96 (0.01) -
@@
Table 7: AUC scores of the algorithms for different instances. The AUC reaches its best score at $1$ and its worst at 00. The values are averages and standard deviations (between brackets) over 16 replications. The “-” entry indicates that an algorithm did not produce results on that instance.
p graph density n glasso RJ-WWA BD-A SS ECM-BAGUS G-MPLBD K-Horseshoe H-BGGM
10 random 10 % 20 0.05 (0.05) 0.09 (0.02) 0.1 (0.02) 0.08 (0.01) 0.18 (0.1) 0.08 (0.02) 0.06 (0.04) 0.88 (0.01)
350 0.05 (0.04) 0.02 (0.01) 0.03 (0.01) 0.03 (0) 0.15 (0.07) 0.01 (0.01) 0.09 (0.06) 0.33 (0.03)
cluster 10 % 20 0.04 (0.03) 0.08 (0.01) 0.09 (0.02) 0.08 (0.01) 0.2 (0.12) 0.07 (0.03) 0.02 (0.02) 0.88 (0.01)
350 0.06 (0.04) 0.02 (0.01) 0.03 (0.01) 0.03 (0) 0.12 (0.05) 0.01 (0.01) 0.07 (0.05) 0.31 (0.02)
100 random 1 % 40 0 (0) 0.05 (0) 0.07 (0) 0.09 (0) 0.16 (0.05) 0.03 (0) 0 (0) -
700 0.01 (0) 0.01 (0) 0.02 (0) 0.03 (0) 0.11 (0.03) 0.01 (0) 0 (0) 0.23 (0)
10 % 40 0.13 (0.1) 0.1 (0.02) 0.12 (0.02) 0.12 (0.01) 0.12 (0.03) 0.04 (0.01) 0.01 (0) -
700 0.17 (0.04) 0.03 (0.01) 0.04 (0.01) 0.05 (0.02) 0.15 (0.04) 0 (0) 0.04 (0) 0.23 (0)
cluster 1 % 40 0 (0) 0.05 (0) 0.07 (0) 0.09 (0) 0.18 (0.04) 0.03 (0) 0 (0) -
700 0 (0) 0.01 (0) 0.02 (0) 0.03 (0) 0.14 (0.04) 0.01 (0) 0 (0) 0.23 (0.01)
10 % 40 0.06 (0.05) 0.08 (0.01) 0.11 (0.01) 0.11 (0.01) 0.12 (0.03) 0.04 (0.01) 0.01 (0) -
700 0.11 (0.02) 0.02 (0) 0.03 (0.01) 0.05 (0.01) 0.12 (0.03) 0 (0) 0.03 (0) 0.23 (0.01)
1000 random 0.1 % 60 0 (0) - - 0.18 (0) 0.29 (0.07) 0.1 (0.01) 0.05 (0) -
1050 0 (0) - - 0.04 (0) 0.26 (0.1) 0 (0) 0.05 (0) -
cluster 0.1 % 60 0 (0) - - 0.18 (0) 0.32 (0.08) 0.1 (0.01) 0.05 (0) -
1050 0 (0) - - 0.04 (0) 0.23 (0.1) 0 (0) 0.05 (0) -
@@
Table 8: $Pr^{-}$ scores (23) of the algorithms for different instances. The $Pr^{-}$ reaches its best score at 00 and its worst at $1$. The values are averages and standard deviations (between brackets) over 16 replications. The “-” entry indicates that an algorithm did not produce results on that instance.
p graph density n glasso RJ-WWA BD-A SS ECM-BAGUS G-MPLBD K-Horseshoe H-BGGM
10 random 10 % 20 0.38 (0.34) 0.38 (0.31) 0.4 (0.31) 0.35 (0.29) 0.31 (0.21) 0.4 (0.33) 0.39 (0.37) 0.9 (0.03)
350 0.87 (0.19) 0.77 (0.23) 0.78 (0.22) 0.56 (0.29) 0.69 (0.24) 0.81 (0.2) 0.85 (0.18) 0.9 (0.11)
cluster 10 % 20 0.45 (0.28) 0.45 (0.23) 0.47 (0.23) 0.4 (0.24) 0.35 (0.19) 0.5 (0.22) 0.45 (0.29) 0.91 (0.02)
350 0.87 (0.15) 0.79 (0.19) 0.8 (0.19) 0.58 (0.23) 0.73 (0.2) 0.82 (0.2) 0.88 (0.17) 0.89 (0.13)
100 random 1 % 40 0.3 (0.15) 0.46 (0.12) 0.5 (0.11) 0.5 (0.11) 0.46 (0.07) 0.47 (0.11) 0.41 (0.12) -
700 0.83 (0.08) 0.84 (0.08) 0.84 (0.08) 0.63 (0.13) 0.82 (0.06) 0.84 (0.07) 0.87 (0.02) 0.9 (0.04)
10 % 40 0.3 (0.11) 0.31 (0.03) 0.32 (0.02) 0.29 (0.02) 0.22 (0.04) 0.22 (0.02) 0.24 (0.03) -
700 0.59 (0.03) 0.73 (0.02) 0.72 (0.02) 0.57 (0.08) 0.58 (0.11) 0.68 (0.02) 0.79 (0.02) 0.76 (0.01)
cluster 1 % 40 0.26 (0.12) 0.56 (0.09) 0.51 (0.08) 0.5 (0.09) 0.46 (0.05) 0.47 (0.08) 0.43 (0.11) -
700 0.81 (0.06) 0.83 (0.05) 0.84 (0.04) 0.61 (0.08) 0.83 (0.04) 0.84 (0.04) 0.81 (0.06) 0.9 (0.03)
10 % 40 0.27 (0.09) 0.33 (0.03) 0.35 (0.03) 0.32 (0.02) 0.24 (0.04) 0.26 (0.03) 0.26 (0.03) -
700 0.59 (0.03) 0.73 (0.03) 0.73 (0.03) 0.59 (0.06) 0.63 (0.05) 0.68 (0.03) 0.8 (0.03) 0.77 (0.02)
1000 random 0.1 % 60 0.24 (0.11) - - 0.61 (0.06) 0.52 (0.07) 0.38 (0.05) 0.62 (0.07) -
1050 0.83 (0.04) - - 0.8 (0.04) 0.76 (0.09) 0.86 (0.03) 0.91 (0.02) -
cluster 0.1 % 60 0.2 (0.15) - - 0.59 (0.09) 0.5 (0.08) 0.36 (0.07) 0.62 (0.08) -
1050 0.8 (0.06) - - 0.78 (0.05) 0.81 (0.08) 0.85 (0.04) 0.91 (0.02) -
@@
Table 9: $Pr^{+}$ scores (22) of the algorithms for different instances. The $Pr^{+}$ reaches its best score at $1$ and its worst at 00. The values are averages and standard deviations (between brackets) over 16 replications. The “-” entry indicates that an algorithm did not produce results on that instance.

### v3

@@ 4 Empirical Comparison
The simulation study in this section contributes to the literature in three ways. First, it includes large-scale instances, i.e., instances with a thousand variables. Until now, Bayesian algorithms are only compared on 250 variables or less, an exception being the K-LRD algorithm by Chandra et al. (2022). Second, we report information on how long it takes algorithms to achieve good results. This information is missing in Bayesian literature. Third, and most importantly, never before has such a variety of Bayesian algorithms been compared in a single simulation study. All our results can be reproduced by the specified scripts on our GitHub page [^1] https://github.com/lucasvogels33/Review-paper-Bayesian-Structure-Learning-in-GGMs. We discuss the simulation setup in Subsection 4.1 and the results in Subsection 4.2.
@@ 4.1 Simulation Setup
Selected algorithms: In Bayesian structure learning literature, new algorithms tend to be compared to frequentist algorithms, or Bayesian algorithms of similar nature, e.g., K-algorithms are compared to K-algorithms, ECM algorithms to ECM algorithms, and so on. We group Bayesian methods into seven categories and show in Table 5 that comparisons between different categories are largely missing in the literature.
@@
Frequentist missing missing 2015 2020 missing 2024 2024 2018 2020
@@
In this section, we complete most of this table by comparing seven Bayesian algorithms (RJ-WWA, BD-A, SS-O, G-MPLBD, K-Horseshoe, ECM-BAGUS, and H-BGGM) and one frequentist algorithm (glasso). For this simulation study, Bayesian algorithms are chosen based on three criteria: i) the algorithm demonstrates outstanding performance in its category, ii) there is publicly accessible working code for the algorithm, and iii) the algorithm is documented in a published article. No regression algorithm is selected because there is no code available. As a prior for the graph we sample the edges from independent and identically distributed Bernoulli distributions. The resulting prior becomes
@@
where $\delta$ is the prior sparsity of the graph. In literature this prior sparsity is often set to either the uninformative value $\delta=0.5$ (Gan et al., 2019) or to a function that decreases with $p$, for example $2/(p-1)$ (Van den Boom et al., 2022a). In this simulation study, we opt for a middle ground and set $\delta=0.2$. Supplementary materials S1 and S2 contain the references and all other settings of the algorithms.
@@
We also measure the computational cost of each algorithm. We first run each Bayesian algorithm until visual inspection shows convergence and denote with $P$ the final edge inclusion matrix. The computational cost $T$ of an algorithm is then defined as the time it takes the algorithm to produce an AUC within 0.01 of its final AUC. That is,
@@
The value of $\delta$ is chosen so that the resulting edge density in the graphs is $10\%$ for $p=10$, $1\%$ and $10\%$ for $p=100$, and $0.1\%$ for $p=1000$. For every generated graph $G^{*}$, the corresponding precision matrix $\mathbf{K}^{*}$ is then sampled from the $G$-Wishart distribution. That is $\mathbf{K}^{*}\sim W_{G^{*}}(3,\mathbf{I})$. For each generated $G^{*}$ and $\mathbf{K}^{*}$, we then sample $n$ observations from the $p-$variate normal distribution with covariance matrix $\mathbf{\Sigma}^{*}=\mathbf{K}^{*{-1}}$ and mean zero. For $p\in\{10,100,1000\}$, each $n$, each graph type, and each density, we repeat the process 16 times which is generally enough to reach low ($<0.025$) standard errors on AUC, $Pr^{+}$ and $Pr^{-}$.
@@ 4.2 Results
The results on the computational costs $T$, AUC, $Pr^{-}$ and $Pr^{+}$ are presented, respectively, in Tables 6, 7, 8, and 9. If there is one conclusion to be drawn from these results, then it is this: the idea that Bayesian algorithms lack the computational advantage of their frequentist counterparts is outdated. Whether with 10, 100, or 1000 variables, one or more Bayesian algorithms can estimate the full posterior distribution or any function thereof at the same time that the glasso algorithm can provide a point estimate. Moreover, there is no evidence of the glasso algorithm outperforming Bayesian algorithms on AUC, $Pr^{+}$, or $Pr^{-}$ either. It underlines the staggering progress Bayesian algorithms have made over the last decade. This progress is illustrated too in Figure 1, which shows the convergence of performance metrics for each algorithm on a large scale instance $(p=1000)$: the AUC, $Pr^{+}$ and $Pr^{-}$ values of the G-MPLBD, SS-O, and K-Horseshoe algorithms converge in approximately the same time as the glasso algorithm.
@@
In terms of the AUC (Table 7), most algorithms perform similarly. Noteworthy is the overall underperformance of the ECM-BAGUS algorithm and the underperformance of the glasso algorithm in medium-sized problems ($p=100$) with a high edge density ($10$%). The $Pr^{-}$ values (Table 8) are comparable across most algorithms too with estimates close to the desired value of zero. Only the ECM-BAGUS and BGGM algorithms perform significantly worse, the latter even assigning in some cases an average edge inclusion probability of $0.88$ to edges that are not in the true graph. Looking at the $Pr^{+}$ values in Table 9, we see more variability between algorithms. For instances with a high number of observations, the K-Horseshoe algorithm performs best. For large-scale graphs, the K-horseshoe algorithm outperforms the others too. The ECM-BAGUS algorithm performs slightly worse than others in most instances. The $Pr^{+}$ values of the BGGM algorithm seem close to optimal but are only marginally more than their $Pr^{-}$ values.
@@
The results also reveal that it is not so much the structure of the graph (Random or Cluster) as the density that influences the performance of algorithms. As expected, more observations lead to better results. Generally, $n=350\log p$ observations are sufficient to yield good results.
@@
Figure 1: AUC (left), $Pr^{+}$ (middle) and $Pr^{-}$ (right) over time. The figures show averages over 16 replications on an instance with $p=1000$, $n=1050$ on the Random graph with density $0.1\%$.
$p$ Graph Density $n$
glasso
RJ-WWA
BD-A
SS-O
ECM-BAGUS
G-MPLBD
K-Horseshoe
H-BGGM
10 Random 10 % 20 2 1 0 0 1 0 0 0
10 % 350 2 0 0 0 1 0 0 0
Cluster 10 % 20 2 1 0 0 1 0 0 0
10 % 350 2 0 0 0 1 0 0 0
100 Random 1 % 40 3 3623 5313 5 84 4 49 -
1 % 700 3 147 1835 19 76 2 3 11
10 % 40 5 2794 9018 10 252 6 32 -
10 % 700 5 746 6093 12 482 8 18 10
Cluster 1 % 40 3 620 4496 13 83 4 11 -
1 % 700 3 159 3013 19 70 1 5 11
10 % 40 4 3489 10368 13 186 5 17 -
10 % 700 3 1569 6465 10 349 6 10 11
1000 Random 0.1 % 60 606 - - 9063 15057 624 24064 -
0.1 % 1050 354 - - 4007 8907 492 12300 -
Cluster 0.1 % 60 672 - - 8510 15710 623 7451 -
0.1 % 1050 366 - - 3823 8224 481 12389 -
@@ 10 Random 10 % 20 2 1 0 0 1 0 0 0
Table 6: Computational cost $T$ (23) in seconds of the algorithms for different instances. The values are averages over 16 replications. The “-” entry indicates that an algorithm did not produce results on that instance.
$p$ Graph Density $n$
glasso
RJ-WWA
BD-A
SS-O
ECM-BAGUS
G-MPLBD
K-Horseshoe
H-BGGM
10 Random 10 % 20 0.7 0.72 0.73 0.73 0.7 0.72 0.7 0.67
10 % 350 0.94 0.95 0.95 0.93 0.91 0.95 0.94 0.93
Cluster 10 % 20 0.83 0.82 0.82 0.81 0.79 0.82 0.81 0.75
10 % 350 0.93 0.93 0.93 0.91 0.92 0.94 0.92 0.92
100 Random 1 % 40 0.83 0.83 0.81 0.84 0.76 0.81 0.84 -
1 % 700 0.96 0.96 0.96 0.94 0.94 0.95 0.96 0.95
10 % 40 0.67 0.7 0.68 0.71 0.59 0.67 0.71 -
10 % 700 0.78 0.92 0.91 0.89 0.8 0.9 0.92 0.87
Cluster 1 % 40 0.83 0.85 0.83 0.83 0.78 0.81 0.85 -
1 % 700 0.95 0.96 0.96 0.94 0.94 0.94 0.95 0.94
10 % 40 0.71 0.77 0.72 0.75 0.61 0.7 0.73 -
10 % 700 0.84 0.94 0.93 0.91 0.84 0.92 0.93 0.87
1000 Random 0.1 % 60 0.86 - - 0.84 0.84 0.72 0.84 -
0.1 % 1050 0.97 - - 0.95 0.96 0.95 0.96 -
Cluster 0.1 % 60 0.84 - - 0.82 0.84 0.7 0.84 -
0.1 % 1050 0.96 - - 0.95 0.95 0.95 0.96 -
@@
Table 7: AUC scores of the algorithms for different instances. The AUC reaches its best score at $1$ and its worst at 00. The values are averages over 16 replications. The “-” entry indicates that an algorithm did not produce results on that instance.
$p$ Graph Density $n$
glasso
RJ-WWA
BD-A
SS-O
ECM-BAGUS
G-MPLBD
K-Horseshoe
H-BGGM
10 Random 10 % 20 0.05 0.09 0.1 0.08 0.18 0.08 0.06 0.88
10 % 350 0.05 0.02 0.03 0.03 0.15 0.01 0.09 0.33
Cluster 10 % 20 0.04 0.08 0.09 0.08 0.2 0.07 0.02 0.88
10 % 350 0.06 0.02 0.03 0.03 0.12 0.01 0.07 0.31
100 Random 1 % 40 0 0.05 0.07 0.09 0.16 0.03 0 -
1 % 700 0.01 0.01 0.02 0.03 0.11 0.01 0 0.23
10 % 40 0.13 0.1 0.12 0.12 0.12 0.04 0.01 -
10 % 700 0.17 0.03 0.04 0.05 0.15 0 0.04 0.23
Cluster 1 % 40 0 0.05 0.07 0.09 0.18 0.03 0 -
1 % 700 0 0.01 0.02 0.03 0.14 0.01 0 0.23
10 % 40 0.06 0.08 0.11 0.11 0.12 0.04 0.01 -
10 % 700 0.11 0.02 0.03 0.05 0.12 0 0.03 0.23
1000 Random 0.1 % 60 0 - - 0.18 0.29 0.1 0.05 -
0.1 % 1050 0 - - 0.04 0.26 0 0.05 -
Cluster 0.1 % 60 0 - - 0.18 0.32 0.1 0.05 -
0.1 % 1050 0 - - 0.04 0.23 0 0.05 -
@@
Table 8: $Pr^{-}$ scores (22) of the algorithms for different instances. The $Pr^{-}$ reaches its best score at 00 and its worst at $1$. The values are averages over 16 replications. The “-” entry indicates that an algorithm did not produce results on that instance.
$p$ Graph Density $n$
glasso
RJ-WWA
BD-A
SS-O
ECM-BAGUS
G-MPLBD
K-Horseshoe
H-BGGM
10 Random 10 % 20 0.38 0.38 0.4 0.35 0.31 0.4 0.39 0.9
10 % 350 0.87 0.77 0.78 0.56 0.69 0.81 0.85 0.9
Cluster 10 % 20 0.45 0.45 0.47 0.4 0.35 0.5 0.45 0.91
10 % 350 0.87 0.79 0.8 0.58 0.73 0.82 0.88 0.89
100 Random 1 % 40 0.3 0.46 0.5 0.5 0.46 0.47 0.41 -
1 % 700 0.83 0.84 0.84 0.63 0.82 0.84 0.87 0.9
10 % 40 0.3 0.31 0.32 0.29 0.22 0.22 0.24 -
10 % 700 0.59 0.73 0.72 0.57 0.58 0.68 0.79 0.76
Cluster 1 % 40 0.26 0.56 0.51 0.5 0.46 0.47 0.43 -
1 % 700 0.81 0.83 0.84 0.61 0.83 0.84 0.81 0.9
10 % 40 0.27 0.33 0.35 0.32 0.24 0.26 0.26 -
10 % 700 0.59 0.73 0.73 0.59 0.63 0.68 0.8 0.77
1000 Random 0.1 % 60 0.24 - - 0.61 0.52 0.38 0.62 -
0.1 % 1050 0.83 - - 0.8 0.76 0.86 0.91 -
Cluster 0.1 % 60 0.2 - - 0.59 0.5 0.36 0.62 -
0.1 % 1050 0.8 - - 0.78 0.81 0.85 0.91 -
@@
Table 9: $Pr^{+}$ scores (21) of the algorithms for different instances. The $Pr^{+}$ reaches its best score at $1$ and its worst at 00. The values are averages over 16 replications. The “-” entry indicates that an algorithm did not produce results on that instance.